On a 5 question multiple-choice test, where each question has 3 answers, what would be the probability of getting at least one question wrong?
Give your answer as a fraction

Answer by Alan3354(69443)

On a 5 question multiple-choice test, where each question has 3 answers, what would be the probability of getting at least one question wrong?
Give your answer as a fraction
------------
Assuming the choices are random, ie, you don't know any of the answers and that all questions are answered:
--
The probability of no wrong answers is 1 in 3^5 = 1 in 243
Any number of wrong answers = 1 - (1/243) = 242/243